# Embedding Similarity Guided License Plate Super Resolution

### v2

@@ 4.2.2 Ablation experiment
To evaluate the impact of embedding dimensionality and distance measures on the LPSR model performance, we conducted a comparative analysis using PSNR and PSNRy. The evaluation spans embedding dimensions of 64, 128, 256, and 512, trained with Euclidean and Manhattan distances. These metrics provide insights into both the fidelity and perceptual quality of the reconstructed images. The results, depicted in Fig. 8, highlight the influence of these factors on model performance on the testing set.
Figure 8: Performance trends of $PSNR_{RGB}$ and $PSNR_{Y}$ across different configurations: embedding distance measures (Euc: Euclidean, Man: Manhattan) and dimensionality $PECL_{d}$ with $d$ the size of the embeddings. The solid lines represent the mean values, while the shaded areas around them indicate the 95% confidence intervals, reflecting the variability or uncertainty of the measurements for each configuration.
From the PSNR curves, the first observation that emerges is that the MSE-based loss lags behind the proposed loss function despite its natural alignment with PSNR, which inherently measures pixel-to-pixel fidelity. This is true independently from the used embedding distance measure. The proposed loss function combines pixel fidelity with embedding similarity, offering a more holistic representation. The adaptive learning mechanism, which dynamically balances the weights of MSE and contrastive losses during training, further boosts model performance. This dual-focus scheme enables the model to effectively capture both fine-grained image details and meaningful feature relationships, resulting in superior super-resolution outcomes.
By comparing among the used distances, the curves highlight clear trends. For PSNR, the Manhattan distance achieves consistently higher values compared to Euclidean across all dimensionalities, with a peak at d=512 where Manhattan reaches approximately 25.0 dB compared to 24.0 dB for Euclidean. Similarly, for PSNRy, Manhattan distance outperforms Euclidean, achieving a maximum of approximately 26.0 dB, compared to 25.0 dB for Euclidean. The gains are evident across all embedding dimensions, particularly in the lower sizes, such as $d=64$ and $d=128$, where Manhattan shows a sharper improvement over the baseline MSE loss function. These results suggest that Manhattan distance offers a more robust representation for feature matching, translating to better perceptual and quantitative image quality during inference. Such behavior can be attributed to its robustness to outliers and its suitability for high-dimensional embedding spaces. Unlike Euclidean distance, Manhattan evaluates differences independently across dimensions, making it less sensitive to extreme values and better aligned with sparse feature representations. This robustness allows Manhattan distance to more effectively capture embedding relationships, leading to improved perceptual and quantitative image quality. The trends also indicate that increasing embedding dimensionality enhances performance for both metrics, with the best results achieved at $d=512$, emphasizing the importance of rich feature representations for accurate embedding similarity measurement.
Figure 9: OCR confidence ($OCR_{\rho}$) trend across different configurations: $PECL_{d}$ with $d$ the size of the embeddings.
By analyzing the accuracy of the detection and readability from the OCR confidence ($OCR_{\rho}$) provided in Fig. 9, we observe that the proposed loss function consistently outperforms the baseline MSE. This comparison highlights the limitations of MSE as a pixel-focused loss function, which primarily optimizes for fidelity at the pixel level but fails to capture feature relationships critical for OCR performance. Additionally, the steady increase in OCR confidence with larger embedding dimensions suggests that richer feature representations enhance OCR accuracy, with the highest performance achieved at d=512. Overall, the training strategy by incorporating embedding similarity with pixel-to-pixel fidelity demonstrates significant advantages over MSE in preserving both perceptual and structural features critical for OCR recognition.
$\text{Contrast}(\text{PECL},MSE)=\frac{\text{PSNR}_{\text{PECL}}-\text{PSNR}_{\text{MSE}}}{\text{PSNR}_{\text{MSE}}}.$ (20)
@@
To evaluate the training dynamics of the proposed PECL, we compare its PSNR performance to the baseline MSE across various embedding sizes (64, 128, 256, and 512). The analysis aims to quantify the relative improvement introduced by PECL using the contrast metric defined in Eq.20. This metric normalizes the difference in PSNR values by the baseline, providing a more interpretable measure of performance gains during training. The results are depicted in Fig.10 and Fig. 11, offering insights into the learning behavior of the proposed PECL with Euclidean and Manhattan distances, respectively. The x-axis, presented on a logarithmic scale, captures training iterations, while the y-axis shows the contrast in PSNR values (in dB). Positive contrast values indicate that the proposed PECL outperforms MSE, while negative values suggest underperformance. The baseline version, MSE, favors high PSNR due to its pixel-wise loss objective. In contrast, the PECL combines MSE with a contrastive loss, which focuses on embedding similarity and perceptual quality, making it less focused on maximizing PSNR directly.
@@
Figure 10: Contrast of PSNR values for PECL with Euclidean distance and embedding sizes of 64, 128, 256, and 512 compared to the MSE during training. The contrast, defined as the relative improvement in PSNR over the MSE, is plotted against the training iterations on a logarithmic scale. Positive contrast values indicate better performance of the PECL models compared to MSE, while negative values indicate worse performance.
@@
Figure 11: Contrast of PSNR values for PECL with Manhattan distance and embedding sizes of 64, 128, 256, and 512 compared to the MSE during training. The contrast, defined as the relative improvement in PSNR over the MSE, is plotted against the training iterations on a logarithmic scale. Positive contrast values indicate better performance of the PECL models compared to MSE, while negative values indicate worse performance.
@@
Upon closer inspection, the contrast curves for Manhattan distance are more centered around the zero axis, with values fluctuating between -0.22 and 0.25. In contrast, the Euclidean distance curves exhibit a wider spread, ranging from -0.6 to 0.13. This difference highlights that, while both distances show a tendency for improvement as training progresses, the Euclidean curves are more influenced by negative contrast values, indicating that Euclidean-based PECL start off with a larger disparity in performance. The more centered Manhattan curves suggest that its impact on the PSNR performance is more stable and balanced throughout training.
@@
During the initial training phase, the contrast values for both Euclidean and Manhattan exhibit fluctuations, reflecting the combined influence of MSE and contrastive loss. As training progresses, the curves stabilize and shift toward positive values, consistently exceeding zero and converging around 0.1. This indicates that PECL, despite improving both MSE and embedding-based similarities, achieves a higher PSNR than MSE. This reflects its ability to produce better visual quality. The positive contrast suggests that the contrastive loss contributes to improving perceptual performance, making the images visually more accurate while still maintaining relatively similar embeddings.
@@
Larger embedding dimensionality exhibits smoother and more stable contrast curves, particularly for Manhattan, suggesting that higher-dimensional embeddings contribute to more robust feature representation and performance stability throughout training. This highlights the role of embedding dimensionality in learning distinctive features, which enhances the model’s ability to generalize. The observed trends emphasize the trade-off managed by PECL, effectively balancing pixel-wise accuracy with perceptual quality to achieve better overall performance compared to the baseline MSE.
@@
Figure 12: Qualitative comparison of license plates under varying acquisition conditions using the proposed PECL with Manhattan and Euclidean distances, compared to the baseline MSE.
@@
The qualitative comparison, as shown in Fig. 12, provides a visual assessment of the performance of the proposed PECL framework with both Manhattan and Euclidean distances, benchmarked against the baseline MSE. The figure presents the high-resolution (HR) reference image alongside the low-resolution (LR) input, downsized by a factor of 8, and the super-resolved (SR) results obtained using MSE, PECL (Man), and PECL (Euc). This layout allows for a direct evaluation of the visual quality achieved by each approach, highlighting the differences in detail preservation, texture sharpness, and perceptual fidelity. As can be seen, the HR images exhibit varying acquisition conditions, including differences in lighting and angles. Additionally, their visual quality appears degraded, with noticeable noise and artifacts. After downsizing them by a factor of 8 to produce the LR inputs, the fine details, such as text and edges, become significantly unclear and, in some cases, unreadable. This highlights the challenge of reconstructing accurate and perceptually meaningful SR outputs.
In the first example, PECL (Man) achieves a PSNR of 29.08 dB and PSNRy of 29.79 dB, outperforming PECL (Euc) and MSE. This improvement is visually apparent, with PECL (Man) generating sharper edges and more detailed textures compared to both MSE and PECL (Euc), which exhibit noticeable smoothing and blurring, particularly in high-frequency regions such as the text. In the second example, PECL (Man) again outperforms both PECL (Euc) and MSE, with a PSNR of 26.14 dB and PSNRy of 26.45 dB. The HR image shows fine details and sharp edges, which are largely preserved in the PECL (Man) reconstruction, while MSE produces a more blurred and less detailed result. This trend is consistent in the third example, where PECL (Man) achieves 25.53 dB in PSNR and 26.78 dB in PSNRy, significantly improving the perceptual quality over PECL (Euc) and MSE, both of which still suffer from substantial blurring and loss of fine details. Across all examples, PECL (Man) consistently restores sharper edges and clearer textures, indicating its ability to recover fine-grained details effectively.
The improved sharpness and perceptual quality in the PECL (Man) results demonstrate the ability of the proposed loss function to better align image features through the embedding consistency enforced by the contrastive loss. In comparison, while PECL (Euc) also improves upon MSE, it still cannot match the sharpness or perceptual quality achieved by PECL (Man), particularly in high-frequency regions. These findings demonstrate that pixel-to-pixel fidelity along with embedding similarity enhances perceptual quality by focusing on feature-level consistency.
@@
Figure 13: Qualitative comparison of HR and SR license plate images, with degradation maps showing noise, blur, and compression artifacts. The maps illustrate the intensity of degradation in both the HR and SR images, with each map providing insights into how the visual quality is affected by the SR model trained with the proposed PECL over MSE.
@@
By analyzing the distortion maps (Noise, Blur, and Compression) in Fig. 13 we can highlight the differences in performance when training the model with MSE compared to PECL with respect to specific degradations. The blur map for PECL demonstrates significantly reduced blur compared to the MSE model, indicating that PECL can restore sharper edges and finer details. In contrast, with MSE, higher blur intensity can be observed around high-frequency regions such as edges. This reflects its tendency to produce over-smoothed outputs. The noise map reveals that PECL maintains moderate and localized noise levels, comparable to MSE. The compression map further demonstrates that PECL introduces fewer compression artifacts, particularly in textured regions, resulting in perceptually superior reconstructions. Compared to the HR image, the proposed method appears to smooth the images, resulting in distribution depicted with the blur maps, regardless of the training loss.

### v3

@@ 4.2.2 Ablation experiment
To validate the efficiency of the proposed PECL over widely used losses, we compare it performances with solely using MSE and MAE. The performances are reported in Table 4. As it can be seen, the proposed PECL, which combines contrastive learning on embeddings with pixel-wise reconstruction, demonstrates consistent superiority over standalone MSE and MAE losses across both the CCPD (large-scale) and PKU (smaller-scale) datasets. On the CCPD dataset, PECL achieves the highest PSNR (25.25 dB) and SSIM (0.8133), outperforming MAE by +0.26 dB and MSE by +1.03 dB in PSNR, while also attaining the best perceptual quality (LPIPS: 0.1031). This indicates that the integration of contrastive loss enhances structural fidelity and pixel-wise accuracy in large-scale settings, likely by leveraging robust embeddings to align high-resolution (HR) and super-resolved (SR) features. Notably, MAE alone outperforms MSE on CCPD, suggesting its outlier robustness benefits dense, diverse data.
On the PKU dataset, PECL again outperforms both baselines, achieving a PSNR of 19.26 dB (+0.56 dB over MSE and +0.65 dB over MAE) and the lowest LPIPS (0.1013). However, MSE slightly surpasses MAE here (PSNR: 18.70 vs. 18.61 dB), likely due to MSE’s quadratic penalty acting as a regularizer in low-data regimes, enforcing stricter pixel-level alignment critical for structured tasks like license plate super-resolution. The contrastive component of PECL mitigates MSE’s tendency to over-smooth textures, as evidenced by improved SSIM (0.7009 vs. 0.6758 for MSE).
Table 4: Performance comparison of the proposed PECL against baseline configurations (MSE, MAE) on the CCPD and PKU datasets. Best performances per dataset are highlighted in bold red.
Loss PSNR PSNRy SSIM SSIMy LPIPS
CCPD
MSE 24.22 ($\pm$2.34) 24.91 ($\pm$2.49) 0.7787 ($\pm$0.06) 0.8143 ($\pm$0.06) 0.1293 ($\pm$0.05)
MAE 24.99 ($\pm$2.49) 25.78 ($\pm$2.64) 0.8118 ($\pm$0.06) 0.8428 ($\pm$0.06) 0.1090 ($\pm$0.04)
PECL 25.25 ($\pm$2.61) 26.12 ($\pm$2.82) 0.8133 ($\pm$0.06) 0.8461 ($\pm$0.06) 0.1031 ($\pm$0.04)
PKU
MSE 18.70 ($\pm$3.76) 18.88 ($\pm$3.86) 0.6758 ($\pm$0.13) 0.6979 ($\pm$0.13) 0.1058 ($\pm$0.04)
MAE 18.61 ($\pm$3.86) 18.78 ($\pm$3.96) 0.6827 ($\pm$0.13) 0.7034 ($\pm$0.13) 0.1023 ($\pm$0.04)
PECL 19.26 ($\pm$3.35) 19.46 ($\pm$3.77) 0.7009 ($\pm$0.12) 0.7222 ($\pm$0.12) 0.1013 ($\pm$0.04)
@@
To evaluate the impact of embedding dimensionality and distance measures on the LPSR model performance, we conducted a comparative analysis using PSNR/PSNRy, SSIM/SSIMy, and LPIPS. The evaluation spans embedding dimensions of 64, 128, 256, and 512, trained with Euclidean and Manhattan distances. These metrics provide insights into both the fidelity and perceptual quality of the reconstructed images. The results, summarized in Table 5 and Table 6, highlight the influence of these factors on model performance.
Table 5: Performance of the proposed PECL w.r.t the embedding distance and dimension on the CCPD dataset [51]. The median ($\pm$ standard deviation) over the testing set is reported. The best performances are highlighted in bold red.
Emb. dim. PSNR PSNRy SSIM SSIMy LPIPS
Manhattan distanc
64 25.10 ($\pm$2.43) 25.92 ($\pm$2.58) 0.8126 ($\pm$0.06) 0.8460 ($\pm$0.06) 0.1075 ($\pm$0.04)
128 25.25 ($\pm$2.61) 26.12 ($\pm$2.82) 0.8133 ($\pm$0.05) 0.8461 ($\pm$0.06) 0.1031 ($\pm$0.03)
256 25.11 ($\pm$2.43) 25.92 ($\pm$2.58) 0.8126 ($\pm$0.06) 0.8460 ($\pm$0.06) 0.1075 ($\pm$0.04)
512 25.12 ($\pm$2.46) 25.93 ($\pm$2.61) 0.8127 ($\pm$0.06) 0.8458 ($\pm$0.06) 0.1065 ($\pm$0.04)
Euclidean distanc
64 24.22 ($\pm$2.38) 24.89 ($\pm$2.52) 0.7511 ($\pm$0.05) 0.7841 ($\pm$0.06) 0.1005 ($\pm$0.04)
128 24.37 ($\pm$2.40) 25.06 ($\pm$2.53) 0.7562 ($\pm$0.07) 0.7875 ($\pm$0.06) 0.0959 ($\pm$0.04)
256 24.54 ($\pm$2.40) 25.24 ($\pm$2.53) 0.7670 ($\pm$0.05) 0.7980 ($\pm$0.05) 0.0926 ($\pm$0.03)
512 24.58 ($\pm$2.45) 25.29 ($\pm$2.58) 0.7715 ($\pm$0.05) 0.8013 ($\pm$0.06) 0.0948 ($\pm$0.03)
Table 6: Performance of the proposed PECL w.r.t the embedding distance and dimension on the PKU dataset [54]. The median ($\pm$ standard deviation) over the testing set is reported. The best performances are highlighted in bold red.
Emb. dim. PSNR PSNRy SSIM SSIMy LPIPS
Manhattan distance
64 18.84 ($\pm$3.83) 19.02 ($\pm$3.94) 0.6821 ($\pm$0.13) 0.7032 ($\pm$0.14) 0.1017 ($\pm$0.04)
128 19.16 ($\pm$3.85) 19.34 ($\pm$3.98) 0.6909 ($\pm$0.12) 0.7124 ($\pm$0.13) 0.1013 ($\pm$0.04)
256 18.61 ($\pm$3.80) 18.79 ($\pm$3.92) 0.6796 ($\pm$0.13) 0.7005 ($\pm$0.13) 0.1006 ($\pm$0.04)
512 18.90 ($\pm$3.89) 19.09 ($\pm$4.00) 0.6905 ($\pm$0.12) 0.7114 ($\pm$0.13) 0.0999 ($\pm$0.04)
Manhattan distance
64 19.26 ($\pm$3.65) 19.46 ($\pm$3.77) 0.7009 ($\pm$0.12) 0.7222 ($\pm$0.12) 0.1013 ($\pm$0.04)
128 19.15 ($\pm$3.65) 19.34 ($\pm$3.75) 0.6933 ($\pm$0.12) 0.7147 ($\pm$0.12) 0.1001 ($\pm$0.04)
256 18.96 ($\pm$3.63) 19.14 ($\pm$3.74) 0.6907 ($\pm$0.12) 0.7119 ($\pm$0.12) 0.0990 ($\pm$0.04)
512 19.08 ($\pm$3.68) 19.28 ($\pm$3.81) 0.6923 ($\pm$0.12) 0.7136 ($\pm$0.12) 0.0987 ($\pm$0.04)
The ablation study on the CCPD dataset, with more than 100k training license plates, demonstrates that the Manhattan-based embedding distance outperforms Euclidean in terms of pixel wise (PSNR) and structural similarity (SSIM) quality metrics, achieving optimal performance at 128 dimensions (PSNR: 25.25 dB, SSIM: 0.8133). This superiority can be attributed to the distance’s robustness to outliers and its ability to preserve fine structural details in high-data regimes, where the large and diverse training set mitigates overfitting risks. The quadratic penalty of MSE, while effective at minimizing large pixel errors, may over-penalize subtle but perceptually acceptable deviations in noisy or variable inputs, such as skewed license plates or lighting variation. This is particularly are prevalent in real-world datasets like CCPD. The contrastive loss component of PECL synergizes with Manhattan distance by learning discriminative embeddings that emphasize structural fidelity, leveraging the dataset’s scale to generalize across diverse samples. The marginal gains beyond 128 dimensions suggest a balance between embedding complexity and reconstruction accuracy, as higher dimensions risk overfitting even in large datasets.
In contrast, the PKU dataset, with 3.5K training license plates, exhibits superior performance with the Euclidean-based embedding distance loss, particularly at 64 dimensions (PSNR: 19.26 dB, SSIM: 0.7009). This reversal highlights the role of dataset size and task-specific regularization. The smaller training set amplifies the Euclidean distance advantages. In particular, its quadratic penalty acts as an implicit regularizer, reducing the risk of overfitting by aggressively minimizing large pixel errors that could otherwise propagate into the embedding space. This is critical for license plate super-resolution, where rigid, high-contrast text demands pixel-perfect alignment to ensure readability. Manhattan distance’s tolerance for moderate errors, while beneficial in large datasets, may fail to enforce the geometric precision required in low-data regimes. Furthermore, the contrastive loss in PECL benefits from Euclidean distance rigor, which stabilizes embedding learning when training samples are limited. The best perceptual quality (LPIPS: 0.0987) at 512 dimensions for Euclidean distance suggests that higher embedding dimensions can enhance perceptual alignment in small datasets, but the minimal gains at lower dimensions (64) underscore the importance of computational efficiency. These results emphasize the task-data interplay, as the Euclidean distance sensitivity to subtle errors aligns with the structured nature of license plates, while its regularization effect compensates for the limited data diversity in PKU dataset.
$\text{Contrast}(\text{PECL},\text{Base})=\frac{\text{PSNR}_{\text{PECL}}-\text{PSNR}_{\text{Base}}}{\text{PSNR}_{\text{Base}}}.$ (20)
@@
To evaluate the training dynamics of the proposed PECL, we compare its PSNR performance to the baseline MSE and MAE losses across various embedding sizes (64, 128, 256, and 512). The analysis aims to quantify the relative improvement introduced by PECL using the contrast metric defined in Eq.20. This metric normalizes the difference in PSNR values by the baseline, providing a more interpretable measure of performance gains during training. The results are depicted in Fig. 10 and Fig. 10, for the contrast with MSE and MAE, respectively. These curves offer insights into the learning behavior of the proposed PECL with Euclidean and Manhattan distances as embedding similarity measures. The x-axis, presented on a logarithmic scale, captures training iterations, while the y-axis shows the contrast in PSNR values (in dB). Positive contrast values indicate that the proposed PECL outperforms the baseline, while negative values suggest underperformance. The baseline versions, MSE and MAE, favors high PSNR due to their pixel-wise loss objective. In contrast, the PECL combines pixel-wise with a contrastive loss, which focuses on embedding similarity and perceptual quality, making it less focused on maximizing PSNR directly.
Refer to caption Refer to caption
@@
Figure 10: PECL and MSE loss relative contrast on the CCPD dataset [51] of PSNR values with embedding sizes of 64, 128, 256, and 512, using Euclidean distance (left) and Manhattan distance (right) during training. The contrast, defined as the relative improvement in PSNR, is plotted against the training iterations on a logarithmic scale. Positive contrast values indicate better performance of the PECL models compared to MSE, while negative values indicate worse performance.
Refer to caption Refer to caption
@@
Figure 11: PECL and MAE loss relative contrast CCPD dataset [51] of PSNR values with embedding sizes of 64, 128, 256, and 512, using Euclidean distance (left) and Manhattan distance (right) during training. The contrast, defined as the relative improvement in PSNR, is plotted against the training iterations on a logarithmic scale. Positive contrast values indicate better performance of the PECL models compared to MSE, while negative values indicate worse performance.
@@
Analyzing the contrast curves obtained on the CCPD dataset in Fig. 10 and Fig. 11, the curves obtained from training the PECL with Manhattan distance are more centered around the zero axis, with values fluctuating between -0.22 and 0.25 when compared to MSE loss, and between -0.3 and 0.1 when compared to MAE loss. In contrast, the Euclidean distance curves exhibit a wider spread, ranging from -0.6 to 0.13 and -0.5 to 0.13, compared to MSE and MAE loss, respectively. This difference highlights that, while both distances show a tendency for improvement as training progresses, the Euclidean curves are more influenced by negative contrast values, indicating that Euclidean-based PECL start off with a larger disparity in performance. The more centered Manhattan curves suggest that its impact on the PSNR performance is more stable and balanced throughout training.
Refer to caption Refer to caption
Figure 12: PECL and MSE loss relative contrast on the PKU dataset [54] of PSNR values with embedding sizes of 64, 128, 256, and 512, using Euclidean distance (left) and Manhattan distance (right) during training. The contrast, defined as the relative improvement in PSNR, is plotted against the training iterations on a logarithmic scale. Positive contrast values indicate better performance of the PECL models compared to MSE, while negative values indicate worse performance.
Refer to caption Refer to caption
Figure 13: PECL and MAE loss relative contrast on the PKU dataset [54] of PSNR values with embedding sizes of 64, 128, 256, and 512, using Euclidean distance (left) and Manhattan distance (right) during training. The contrast, defined as the relative improvement in PSNR, is plotted against the training iterations on a logarithmic scale. Positive contrast values indicate better performance of the PECL models compared to MSE, while negative values indicate worse performance.
@@
Similarly, the curves obtained on the PKU dataset, depicted in Fig. 12 and Fig. 13 showcase the same behavior. During the initial training phase, the contrast values for both Euclidean and Manhattan exhibit fluctuations, reflecting the combined influence of MSE and contrastive loss. As training progresses, the curves stabilize and shift toward positive values, consistently exceeding zero and converging around 0.1 and 0.2 on the CCPD and PKu datset, respectively. This indicates that PECL, despite improving both pixel-wise and embedding-based similarities, achieves a higher PSNR than MSE or MAE. This reflects its ability to produce better visual quality. The positive contrast suggests that the contrastive loss contributes to improving perceptual performance, making the images visually more accurate while still maintaining relatively similar embeddings.
@@
Larger embedding dimensionality exhibits smoother and more stable contrast curves, regardless of which baseline loss is compared to. Particularly for Manhattan, suggesting that higher-dimensional embeddings contribute to more robust feature representation and performance stability throughout training. This highlights the role of embedding dimensionality in learning distinctive features, which enhances the model’s ability to generalize. The observed trends emphasize the trade-off managed by PECL, effectively balancing pixel-wise accuracy with perceptual quality to achieve better overall performance compared to the baseline MSE.
@@
Figure 14: Qualitative comparison of license plates from the CCPD dataset [51] under varying acquisition conditions using the proposed PECL with Manhattan and Euclidean distances, compared to the baseline MSE and MAE losses.
Figure 15: Qualitative comparison of license plates from the PKU dataset [54] under varying acquisition conditions using the proposed PECL with Manhattan and Euclidean distances, compared to the baseline MSE and MAE losses.
@@
The qualitative comparison, as shown in Fig. 14 and Fig. 15, provides a visual assessment of the performance of the proposed PECL framework with both Manhattan and Euclidean distances, benchmarked against the baseline MSE and MAE losses. The figures present the high-resolution (HR) reference image alongside the low-resolution (LR) input, downsized by a factor of 8, and the super-resolved (SR) results obtained using MSE, MAE, PECL (Man), and PECL (Euc). This layout allows for a direct evaluation of the visual quality achieved by each approach, highlighting the differences in detail preservation, texture sharpness, and perceptual fidelity. As can be seen, the HR images exhibit varying acquisition conditions, including differences in lighting and angles. Additionally, their visual quality appears degraded, with noticeable noise and artifacts. After downsizing them by a factor of 8 to produce the LR inputs, the fine details, such as text and edges, become significantly unclear and, in some cases, unreadable. This highlights the challenge of reconstructing accurate and perceptually meaningful SR outputs.
From the CCPD dataset, PECL (Man) clearly outperforms the alternatives: in the first example, it achieves a PSNR of 29.08 dB and PSNRy of 29.79 dB, producing significantly sharper edges and more detailed textures than PECL (Euc), MSE, or MAE, which tend to smooth or blur high-frequency details such as text. In subsequent examples, PECL (Man) maintains this advantage, with PSNR values of 26.14 dB and 25.53 dB (with corresponding PSNRy improvements), thereby preserving the fine structural details evident in the high-resolution references. Conversely, for samples drawn from the PKU dataset, PECL (Euc) consistently outperforms the other losses, attaining a PSNR of 21.17 dB and PSNRy of 21.48 dB in the first sample by delivering visually sharper reconstructions and more refined textures compared to MSE, MAE, and even PECL (Man). These observations collectively underscore the benefit of integrating embedding consistency through contrastive loss: while PECL (Euc) enhances structural fidelity under certain conditions, the embedding alignment enforced by PECL (Man) appears particularly adept at recovering fine-grained details and ensuring perceptual quality, especially in high-frequency regions.
It is also important to highlight the quality differences between the datasets. The CCPD dataset is large-scale, encompassing a wide variety of imaging conditions, which facilitates a more robust evaluation of super-resolution methods. On the other hand, the PKU dataset is small-scale, characterized by a limited number of images and generally lower inherent quality. This distinction emphasizes the adaptability of the proposed approach, where PECL (Man) proves particularly effective on the more varied and higher-quality CCPD dataset, while PECL (Euc) shows relative advantages on the smaller-scale, more challenging PKU dataset.
@@
Figure 16: Qualitative comparison of HR and SR license plate images, with degradation maps showing noise, blur, and compression artifacts. (top) CCPD dataset [51] and (bottom) PKU dataset [54]. The maps illustrate the intensity of degradation in both the HR and SR images, with each map providing insights into how the visual quality is affected by the SR model trained with the proposed PECL over MSE and MAE losses.
@@
By analyzing the distortion maps (Noise, Blur, and Compression) in Fig. 16 we can highlight the differences in performance when training the model with MSE or MAE compared to PECL, with respect to specific degradations. The blur map for PECL demonstrates significantly reduced blur, regardless on which dataset, indicating that PECL can restore sharper edges and finer details. In contrast, with MSE and MAE, higher blur intensity can be observed around high-frequency regions such as edges. This reflects its tendency to produce over-smoothed outputs. The noise map reveals that PECL maintains moderate and localized noise levels, comparable to MSE. The compression map further demonstrates that PECL introduces fewer compression artifacts, particularly in textured regions, resulting in perceptually superior reconstructions. Compared to the HR image, the proposed method appears to smooth the images, resulting in distribution depicted with the blur maps, regardless of the training loss.
@@
Embedding space visualization: We conduct a feature visualization using t-SNE visualizations of feature embeddings among Euclidean and Manhattan distances as embedding similarity measures.
The t-SNE visualizations of feature embeddings on the CCPD dataset, shown in Fig.17, indicate that both distance metrics produce compact and well-clustered embeddings in the latent space. Notably, Manhattan distance results in slightly tighter clustering compared to Euclidean distance, suggesting that it enforces a stronger alignment between HR and SR embeddings. However, the visualization on the PKU dataset, shown in Fig.18, exhibits a different pattern. The embeddings appear more distinct, with well-separated clusters likely influenced by the varying characteristics of license plate images, particularly differences in background color. This distribution suggests that factors beyond the choice of distance metric contribute to the structure of the latent space in this dataset. Besides, the embedding distribution on the PKU dataset highlights it challenging nature, yet the model is able to generate SR images with similar embeddings to the HR one for several data points.
Refer to caption Refer to caption
Figure 17: t-SNE visualization of feature embeddings for high-resolution and super-resolved images, with respect to the embedding distance on the CCPD dataset [51]. (Left) Euclidean distance and (right) Manhattan distance.
Refer to caption Refer to caption
Figure 18: t-SNE visualization of feature embeddings for high-resolution and super-resolved images, with respect to the embedding distance on the PKU dataset [54]. (Left) Euclidean distance and (right) Manhattan distance.
4.2.3 Applicability to $\times 4$ Super-Resolution
While the primary focus of this work is on extreme $\times 8$ super-resolution of license plate images, we also include an additional comparison at $\times 4$ to align with recent literature and assess the generalization ability of the proposed method.
We adapted our model to operate under a $\times 4$ upscaling setting and compared its performance to four recent and competitive methods: SwinIR [28], DiT-SR [6], DiffTSR [58], and SeeSR [50]. Performance is studied on the CCPD dataset [51] because of its large size and the availability of labels. These models were selected based on their strong relevance and performance in text and general image super-resolution tasks.
Quantitative results are presented in Table 7, focusing on visual fidelity (PSNR, SSIM) and perceptual similarity (LPIPS). Despite being optimized for the more challenging $\times 8$ scale, our model demonstrates strong performance at $\times 4$, indicating robust generalization. Among the compared methods, SeeSR performs the weakest, showing the lowest PSNR and SSIM scores and the highest LPIPS, which reflects poor reconstruction and weak perceptual quality. DiT-SR performs better but still lags behind the other approaches. SwinIR offers notable improvements in both PSNR and SSIM, along with relatively low LPIPS, highlighting its strength in preserving both details and structure. DiffTSR further improves upon SwinIR, achieving the second-best results across all metrics, particularly with a lower LPIPS score (0.0581). Our proposed model outperforms all competitors, achieving the highest PSNR (31.45), SSIM (0.8590), and the lowest LPIPS (0.0436), demonstrating its effectiveness in producing visually accurate and perceptually faithful images under moderate upscaling.
Beyond visual quality, Table 8 evaluates the impact of super-resolution on license plate recognition. The results reveal consistent trends. SeeSR again performs the worst, with poor accuracy and high error rates, indicating difficulty in preserving readable text. DiT-SR shows moderate improvements but remains below SwinIR and DiffTSR. SwinIR delivers strong recognition results, particularly in exact match accuracy and character-level metrics, confirming its reliability. DiffTSR improves further, ranking second in most recognition metrics, including word error rate and exact match.
Our model achieves the best performance overall, with the highest exact match accuracy (69.60%), the lowest character error rate (9.80%), and the top scores in precision, recall, and F1-score. These results highlight the model’s ability to restore sharp and readable text details, which is crucial for accurate license plate recognition. In summary, even though it is primarily designed for $\times 8$ super-resolution, our method proves highly effective and competitive at $\times 4$, both in terms of visual quality and practical downstream recognition.
Table 7: Performance comparison of the proposed model with recent SR methods on the CCPD dataset [51] under $\times 4$. The median ($\pm$ standard deviation) over the testing set is reported. The best and second-best performances are respectively highlighted in bold red and bold blue.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SwinIR [28] 30.94 ($\pm$ 1.29) 0.8396 ($\pm$ 0.094) 0.0792 ($\pm$ 0.027)
DiT-SR [6] 29.73 ($\pm$ 0.94) 0.5825 ($\pm$ 0.063) 0.1546 ($\pm$ 0.048)
SeeSR [50] 28.26 ($\pm$ 0.57) 0.5251 ($\pm$ 0.092) 0.2032 ($\pm$ 0.080)
DiffTSR [58] 31.26 ($\pm$ 1.30) 0.8412 ($\pm$ 0.082) 0.0581 ($\pm$ 0.010)
Ours (×4) 31.45 ($\pm$ 1.54) 0.8590 ($\pm$ 0.102) 0.0436 ($\pm$ 0.015)
Table 8: Text recognition performance on super-resolved license plates from the CCPD dataset [51] under $\times 4$. The best and second-best results are highlighted in bold red and bold blue, respectively.
[…]
SwinIR [28] 67.10% 88.82% 11.18% 32.90% 83.27% 83.35% 83.21%
DiT-SR [6] 56.80% 84.57% 15.43% 43.20% 76.66% 76.75% 76.57%
SeeSR [50] 28.99% 49.14% 59.86% 60.01% 42.23% 42.35% 41.10%
DiffTSR [58] 69.09% 89.95% 12.05% 31.80% 82.12% 82.10% 82.00%
Ours (×4) 69.60% 90.20% 9.80% 30.40% 85.52% 85.54% 85.45%